a box for shipping roofing nails must have a volume of 84 cubic feet. if the box must be 3 feet wide and its height must be 3 feet less then its length,what should the dimensions of the box be?
Answer by RAY100(1637)

Volume =84 = L*W*H = L*(L-3)*(3)= 3L^2 -9L
.
3L^2 -9L -84 =0,,,,,divide by 3
.
L^2 -3L-28 = 0
.
(L-7)(l+4) =0
.
L=7,,-4,,,not realistic
.
.
check, L=7,,,,H=7-3=4,,,W=3,,,,,Vol = 7*4*3 =84,,,ok
.